OpenStudy (anonymous):

How can one fifth x − 2 = one third x + 8 be set up as a system of equations? Answer options: a. 5y − 5x = −10 3y − 3x = 24 b. 5y − 5x = −10 3y + 3x = 24 c. 5y + x = −10 3y + x = 24 d. 5y − x = −10 3y − x = 24 Thanks in advance!

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{5}x-2=\cfrac{1}{3}x+8?\)

OpenStudy (anonymous):

@jdoe0001 correct.

OpenStudy (mrnood):

this can be solved without using 'simultaneous equations' since there is only one variable x. However you can take each side of the equation and set up as a separate equation y= x/5 -2 y= x/3 +8 Just rearrange THOSE equations until they give you one of the answers

OpenStudy (jdoe0001):

\(\large { \begin{array}{cccllll} \cfrac{1}{5}x-2&=&\cfrac{1}{3}x+8\\ y&=&y \\\hline\\ y=\cfrac{1}{5}x-2&&y=\cfrac{1}{3}x+8 \end{array}\implies \begin{cases} y=\cfrac{1}{5}x-2\\ y=\cfrac{1}{3}x+8 \end{cases} }\) as MrNood indicated

OpenStudy (jdoe0001):

now, neither of those look like any of the choices but you can simplify them some by 1st one -- multiply both sides by 5 2nd one -- multiply both sides by 3

OpenStudy (mrnood):

try doing what @jdoe0001 suggested above multiply those equaitons and the answer will be clear

OpenStudy (mrnood):

y= x/5 -2 multiply this by 5 both sides y= x/3 +8 multiply this by 3 both sides

OpenStudy (perl):

This is not an efficient way to solve for x, x − 2 = 1/3 (x + 8) treat each side as separate functions y= x/5 - 2 y= x/3 +8 now clear the fractions

OpenStudy (mrnood):

This is exactly as 2 people have posted above and btw: you have a typo it's x/5-2 = x/3 + 8 as confirmed by OP your brackets are in the wrong place on right hand side. As noted above there is only one variable so this can be solved without simultaneous equations

OpenStudy (anonymous):

@LelandP

OpenStudy (mrnood):

@Isherron @LelandP No need for other shout outs - the answer is clearly give by no less than THREE competent people above Just DO the multiplication - the answer is there...